# Strip R-CNN: Large Strip Convolution for Remote Sensing Object Detection

### v3

@@ 4.5 Visual Analysis
We also present detection results and Eigen-CAM [50] visualizations in Fig. 8. From these images, it is evident that for high aspect ratio objects, previous methods such as Oriented-RCNN and LSKNet exhibit issues like missed detections and significant localization errors. In contrast, our method can successfully detect the high aspect ratio objects. The Eigen-CAM visualizations also show strong activations of our method for these objects, further validating that our approach can alleviate the challenges associated with detecting high aspect ratio objects.
@@ 5 Conclusions
In this paper, we alleviate the challenge of detecting slender objects in remote sensing scenarios by leveraging large strip convolutions to better extract features and improve localization of such objects. Based on large strip convolutions, we propose the simple yet highly effective Strip R-CNN. Extensive experiments demonstrate that our method exhibits strong generalization capability and achieve state-of-the-art performance on several remote sensing benchmarks. We hope this research could facilitate the development of object detection in the remote sensing field.
@@ 8 FAIR1M benchmark results
9 DOTA-v1.5 benchmark results
Fine-grained category result comparisons with state-of-the-art methods on the DOTA-v1.5 dataset are given in Tab. 14

### v4

@@ 4.4 Visual Analysis
We also present detection results and Eigen-CAM [48] visualizations in Fig. 9. We can see that for high aspect ratio objects, previous methods such as Oriented-RCNN and LSKNet have issues like missed detections and significant localization errors. In contrast, our method can successfully detect the high aspect ratio objects. The Eigen-CAM visualizations also show strong activations of our method for these objects, validating the effectiveness of our approach.
@@ 5 Conclusions
In this paper, we alleviate the challenge of detecting slender objects in remote sensing scenarios by leveraging large strip convolutions to better extract features and improve localization of such objects. Based on large strip convolutions, we propose the simple yet highly effective Strip R-CNN. Extensive experiments demonstrate that our method exhibits strong generalization capability and achieves state-of-the-art performance on several remote sensing benchmarks. We hope this research could facilitate the development of object detection in the remote sensing field.
@@ 8 FAIR1M benchmark results
9 DOTA benchmark results
Fine-grained category result comparisons with state-of-the-art methods on the DOTA-v1.0 and DOTA-v1.5 dataset are given in Tab. 16 and  Tab. LABEL:dota1.0_sup
Single-Scale
COBB [63] IN 76.52 41.9M - - - - - - - - - - - - - - - -
Multi-Scale
R3Det [68] IN 76.47 41.9M 336G 89.80 83.77 48.11 66.77 78.76 83.27 87.84 90.82 85.38 85.51 65.57 62.68 67.53 78.56 72.62
S2ANet [19] IN 79.42 - - 88.89 83.60 57.74 81.95 79.94 83.19 89.11 90.78 84.87 87.81 70.30 68.25 78.30 77.01 69.58
R3Det-GWD [71] IN 80.23 41.9M 336G 89.66 84.99 59.26 82.19 78.97 84.83 87.70 90.21 86.54 86.85 73.47 67.77 76.92 79.22 74.92
R3Det-KLD [72] IN 80.63 41.9M 336G 89.92 85.13 59.19 81.33 78.82 84.38 87.50 89.80 87.33 87.00 72.57 71.35 77.12 79.34 78.68
PKINet-S [1] IN 81.06 30.8M 190G 89.02 86.73 58.95 81.20 80.41 84.94 88.10 90.88 86.60 87.28 67.10 74.81 78.18 81.91 70.62